OpenStudy (anonymous):

Suppose you choose a marble from a bag containing 4 red marbles, 2 white marbles, and 3 blue marbles. You return the first marble to the bag and then choose again. Find P(red and blue).

OpenStudy (anonymous):

you mean first red and blue, or just one red and on blue order not important?

OpenStudy (anonymous):

if it is first red and then blue, should be \(\frac{4}{9}\times \frac{3}{9}\) i.e. multiply the probabilities together

OpenStudy (anonymous):

A. 4/3 B. 2/3 C. 7/9 D. 4/27

OpenStudy (anonymous):

ASHWORTH lol c(:

OpenStudy (anonymous):

(D)